OpenStudy (anonymous):

A rectangular prism has length thirteen meters, width six meters, and volume three hundred ninety cubic meters. What is the height of the prism? A. four meters B. five meters C. nineteen meters D. seventy-eight meters

OpenStudy (jdoe0001):

volume of a rectangular prism = length x width x height length = 13m widht = 6m height = h volume = 390 so \(\begin{matrix} volume &=& length& \times &width& \times &height\\ 390&=&13& \times & 6 & \times & h \end{matrix}\) solve for "h"

OpenStudy (anonymous):

o-o thanks